OpenStudy (anonymous):

Identify the vertex for the graph of y = −2x2 + 8x − 1. a. (−2, 7) b. (2, 7) c. (2, −9) d. (−2, −9)

OpenStudy (anonymous):

first coordinate of the vertex is \(-\frac{b}{2a}\) which in your case is \[-\frac{8}{2\times -2}=2\]

OpenStudy (igreen):

First find the axis of symmetry. \(-\dfrac{b}{2a}\)

OpenStudy (anonymous):

second coordinate of the vertex is what you get for \(y\) when you replace \(x\) by \(2\)

OpenStudy (igreen):

^

OpenStudy (igreen):

\(y = −2x^2 + 8x − 1\) \(y = −2(2)^2 + 8(2) − 1\) \(y = −2(4) + 16 − 1\) \(y = -8 + 16 - 1\)

OpenStudy (igreen):

Finish it @Shadowgirl15

OpenStudy (anonymous):

IDK how

OpenStudy (kainui):

 The vertex is the point where if you go left and right by the same amount, you'll end up at the same height, y. So if you have \[\Large y=ax^2+bx+c\] then if we say our vertex point is v then we can travel left or right by x and we will have the same y value, so we can plug this in: \[\Large y=a(v+x)^2+b(v+x)+c=a(v-x)^2+b(v-x)+c\] So now we can solve for v, do a little algebra: \[\Large a[(v+x)^2-(v-x)^2]=b[(v-x)-(v+x)]\]\[\Large a4vx=-2bx\]\[\Large v=\frac{-b}{2a}\] So that's the point, derived with algebra.

OpenStudy (anonymous):

ok but i still dont know the answer

OpenStudy (kainui):

I don't understand. The answer is given by several people before me.

OpenStudy (anonymous):

so is it b?